The length of a rectangular banner is 2 feet longer than its width. If the area is 80 square feet, find the dimensions.

Answer by checkley79(3341)

AREA=LW
L=W+2
80=(W+2)W
80=W^2+2
W^2+2-80=0
(W+10)(W-8)=0
W-8=0
W=8 FEET FOR THE WIDTH.
L=8+2=10 FEET FOR THE LENGTH.
PROOF:
80=10*8
80=80